OpenStudy (anonymous):

The two triangles below are similar. What is the similarity ratio of ΔABC to ΔXYZ? Triangle XYZ is shown with XZ measuring 10 and YZ measuring 30. Triangle ABC is shown with side AC measuring 15. 1:2 2:1 3:2 2:3

OpenStudy (abmon98):

AB/XY=BC/YZ=AC/XZ AB/XY=BC/YZ=15/10 3:2 which is the similarity Ratio.